Name: aux pobles Show all relevant work for full credit. Ch.3 08/26/2019 Homework #3 - Math 227 Use the following data to ans can someone help with step by step answers to these problems
Answer #1
Sr. no X (X - X̄)²
1 1.4 43.12
2 2 35.60
3 2.5 29.88
4 3 24.67
5 3.3 21.78
6 3.7 18.20
7 3.7 18.204
8 4.2 14.188
9 4.4 12.721
10 4.6 11.334
11 4.8 10.028
12 6.7 1.604
13 6.8 1.361
14 7.1 0.751
15 7.3 0.444
16 8.8 0.694
17 9.1 1.284
18 9.4 2.054
19 9.4 2.054
20 10.2 4.988
21 10.4 5.921
22 10.6 6.934
23 11 9.201
24 11.7 13.938
25 12.6 21.468
26 12.9 24.3378
27 13.3 28.4444
28 14.2 38.8544
29 14.4 41.3878
30 15.5 56.7511
X (X - X̄)²
total sum 239 502.21
n 30 30

a)mean =    ΣX/n =    239.000   /   30   =   7.9667

b)sample std dev =   √ [ Σ(X - X̄)²/(n-1)] =   √   (502.2067/29)   =       4.1614

c)Median=0.5(n+1)th value =    15.5th   value of sorted data
=   8.05  

d)mode= highest frequency data =    3.7 , 9.4

e)range=max-min =    15.5   -   1.4   =   14.1

f) sample variance =    Σ(X - X̄)²/(n-1)=   502.2067   /   29   =   17.317

g) mid range=(max + min)/2= (   15.5   +   1.4   ) /2 =    8.45

h) D6 = 0.6*(n+1)th value = 18.6th value = 9.4

i) percentile of CA = number of values below CA/n = 23/30=0.7667

so, percentile = 76.76th

j) quartile , Q1 = 0.25(n+1)th value=   7.75th   value of sorted data
=   4.075  
      
Quartile , Q3 = 0.75(n+1)th value=   23.25th   value of sorted data
=   11.175  
      
IQR = Q3-Q1 =    7.1  
      
1.5IQR =    10.65  
      
lower bound=Q1-1.5IQR=   -6.58  
      
upper bound=Q3+1.5IQR=   21.83  
      
outlier =values outside lower bound and upper bound      
total outlier below lower bound=   0  
total outlier above upper bound=   0  
total outlier =    0